# Are Large Reasoning Models Interruptible?

### v1

@@ 4 How do models behave under time-constraint interruptions?
In terms of accuracy, under soft interrupts, models perform comparable to the full thinking mode across the interrupt positions on easy tasks (see Figure E.1). However, on more challenging tasks such as AIME and LiveCodeBench-v6, we observe cases of panic: models prematurely terminate their reasoning after receiving the speedup instruction, leading to up to 30% accuracy drops. As shown in Figure 1(b), GPT-OSS and Qwen3 sometimes abandon their reasoning within a few tokens after the update, with up to 80% of performance loss attributable to this panic behavior. A qualitative example is provided in subsection C.2.
@@ 5 How do models behave under update-driven interruptions?
In addition to interrupting without new information for efficiency or to obtain an immediate answer, users may also want to interrupt the model with an updated context, or with some new information for the model to incorporate into their solution. We interrupt models at different points in their reasoning trace with new information for their downstream task, as discussed in subsection 3.2. The results are shown in Figure 4 (w/o prompt guidance), where we see that updates lead to dramatic drops in performance, particularly for late-stage interruptions, where models are unable to continue their thinking traces and recover from updates to the underlying problem specification. One of the reasons that we find for this drop is a phenomenon that we call “self-doubt”: an example is provided in Figure 5 (red). Here, models are prone to doubting whether the update is correct and continue with their original thinking process without taking into account the new updated information, even when warned that updates are expected in the initial system prompt (see Appendix B). To remedy this behavior, we introduce “prompt guidance”, a short postfix string appended after the update tag in the model’s voice, stating that the update is correct and verified by the user (Figure 5, green). While prompt guidance substantially improves performance on easier tasks such as GSM8K and MATH-500, AIME tasks still show considerable room for improvement. For coding tasks, models remain particularly fragile when handling mid-reasoning updates, especially when the update involves changes to starter code.
@@
In summary, while models struggle to incorporate mid-reasoning updates without guidance, often exhibiting “self-doubt” and ignoring new information, especially during late-stage interruptions, carefully designed prompt strategies can substantially enhance their adaptability.
@@ 6.2 Assistant-turn vs User-turn Interruption
We also conducted experiments on update-driven interruptions. As shown in Figure 8, we evaluate interventions at both the user-turn and assistant-turn levels. User-turn interventions perform slightly worse than assistant-turn updates when using our prompt guidance, although both outperform the unguided baseline. The difference is most pronounced for Magistral models on GSM8K and MATH-500. On AIME, prompt guidance also provides a clear advantage over naïve user-turn interruptions.
It is worth noting that Qwen3 and Magistral models often struggle with consistent formatting. We therefore relaxed formatting constraints, counting an answer as correct if the final result matched, regardless of output structure. Even with this adjustment, their overall performance remained relatively low. For these reasons, we adopted the assistant-turn interruption setup for our main experiments. Future work may further investigate user-turn interruptions, which could be more natural in interactive deployments.

### v3

@@ 4 How do models behave under time-constraint interruptions?
In terms of accuracy, under soft interrupts, models perform comparable to the full thinking mode across the interrupt positions on easy tasks (see Figure E.1). However, on more challenging tasks such as AIME and LiveCodeBench-v6, we observe cases of panic: models prematurely terminate their reasoning after receiving the speedup instruction; more concretely, panic behavior is defined as the model closing its thinking after using less than 1% of its left context limit after the soft interrupt. This pathology results in up to 30% accuracy drops. As shown in Figure 1(b), GPT-OSS and Qwen3 sometimes abandon their reasoning within a few tokens after the update, with up to 80% of performance loss attributable to this panic behavior. A qualitative example is provided in subsection C.2.
@@ 5 How do models behave under update-driven interruptions?
In addition to interrupting without new information for efficiency or to obtain an immediate answer, users may also want to interrupt the model with an updated context, or with some new information for the model to incorporate into their solution. We interrupt models at different points in their reasoning trace with new information for their downstream task, as discussed in subsection 3.2. The results are shown in Figure 4 (w/o prompt guidance), where we see that updates lead to drops in performance, particularly for late-stage interruptions, where models are unable to continue their thinking traces and recover from updates to the underlying problem specification. One of the reasons that we find for this drop is a phenomenon that we call “self-doubt": an example is provided in Figure 5 (red). Here, models are prone to doubting whether the update is correct and continue with their original thinking process without taking into account the new updated information, even when warned that updates are expected in the initial system prompt (see Appendix B). The doubt rates across models and datasets are shown in Figure 1(b, right): (1) the pathology is more pronounced in math tasks, and (2) GPT-OSS exhibits minimal doubt behavior compared to the other models. More concretely, we classify doubt behavior across failure cases using an LLM-based classifier (GPT-5), with the update and the truncated reasoning trace after the interrupt (potentially containing doubt behavior) as inputs to the model.
To remedy this behavior, we introduce “prompt guidance”, a short postfix string appended after the update tag in the model’s voice, stating that the update is correct and verified by the user (Figure 5, green). While prompt guidance substantially improves performance on easier tasks such as GSM8K and MATH-500, AIME tasks still show considerable room for improvement. For coding tasks, models, specifically Qwen3, remain particularly fragile when handling mid-reasoning updates.
@@
In summary, while models struggle to incorporate mid-reasoning updates without guidance, often exhibiting “self-doubt” and ignoring new information, especially during late-stage interruptions, carefully designed prompt strategies can improve their adaptability.
@@ 6.2 Assistant-turn vs User-turn Interruption
We also conducted experiments on update-driven interruptions by inserting a new user turn mid-reasoning. As shown in Figure 8, we evaluate interventions at both the user-turn and assistant-turn levels. User-turn interventions perform slightly worse than assistant-turn updates when using our prompt guidance, although both outperform the unguided baseline. The difference is most pronounced for the Magistral model on GSM8K and MATH-500. On AIME, prompt guidance also provides a clear advantage over naive user-turn interruptions.
It is worth noting that Qwen3 and Magistral models often struggle with consistent formatting. We therefore relaxed formatting constraints, counting an answer as correct if the final result matched, regardless of the output structure. Even with this adjustment, their overall performance remained relatively low. For these reasons, we adopted the assistant-turn interruption setup for our main experiments. Future work may further investigate user-turn interruptions, which could be more natural in interactive deployments.